# Revisiting electron-capture decay for Galactic cosmic-ray data

### v1

@@ 2.1 From differential density to modulated fluxes (without EC)
$\begin{split}-\vec{\nabla}\left[D(E)\vec{\nabla}N_{j}-\vec{V}_{c}N_{i}\right]+\frac{\partial}{\partial E}\left[b_{\rm tot}(E)N_{i}-(v/c)^{2}K_{pp}\frac{\partial N_{i}}{\partial E}\right]\\ +\left(\Gamma^{\rm inel}_{i}+\Gamma^{\beta}_{i}\right)N_{i}=q_{i}+\sum_{j>i}\Gamma^{\rm prod}_{j\to i}N_{j}+\sum_{k>i}\Gamma^{\beta}_{k\to i}N_{k}\,.\end{split}$ (1)
In the first line, from left to right, the transport terms are: spatial diffusion $D$ (CR scattering off the magnetic turbulence), convective transport $V_{c}$, energy losses (ionisation, Coulomb, and adiabatic) and gains (momentum diffusion $K_{pp}$). In the second line, the CR destruction (left-hand side) and production from heavier species (right-hand side) involve nuclear cross-sections
$\Gamma^{\rm inel}_{i}=\sum_{t\,\in\,\rm ISM}\,n_{t}\,v_{i}\,\sigma_{i+t}^{\rm inel}\quad{\rm and}\quad\displaystyle\Gamma^{\rm prod}_{j\to i}=\sum_{t\,\in\,\rm ISM}\,n_{t}\,v_{i}\,\sigma_{j+t\to i}^{\rm prod}\,$, (2)
and $\beta$ decay (half-life $t_{1/2}$) as a sink (decay of $i$) or a source (decay of CR $k$ into $i$),
@@
Finally, the injection rate $q_{i}$ describes the primary production.
This transport equation can be solved by different approaches and associated codes: numerically after discretisation on spatial and energy coordinates – e.g., DRAGON (Evoli et al., 2018), GALPROP (Porter et al., 2022) and PICARD (Kissmann, 2014) –, via stochastic differential equation based Monte Carlo simulations (e.g., CRPROPA, Merten and Aerdker 2025), or from semi-analytical solutions in simplified geometries based on Green functions, series expansion, etc. (e.g., the LBM and 1D-DM in USINE [^1] https://dmaurin.gitlab.io/USINE/, Maurin 2020). Once $N$ is calculated for all CR isotopes, the (nearly isotropic) CR interstellar (IS) flux is given by $\psi^{\rm IS}=N(v/4\pi)$. To compare the modelled fluxes to the measured ones, Solar modulated top-of-atmosphere (TOA) fluxes are calculated from IS ones, using the force-field approximation (Gleeson and Axford, 1967, 1968). The energy is transformed according to
@@
with $Z$ and $A$ the atomic and mass number of the species, $\phi_{\rm FF}$ the solar modulation potential varying from $\sim 500$ MV (Solar minimum) to 1200 MV (Solar maximum). The flux at the top of the atmosphere is related to the interstellar flux by $\psi^{\rm TOA}(E^{\rm TOA})=\psi^{\rm IS}(E^{\rm IS})\times(p^{\rm TOA}/p^{\rm IS})^{2}$. In the above equations, $p$ is the momentum, $E$ is the total energy, and $E_{\rm k/n}=E_{\rm k}/A$ is the kinetic energy per nucleon, with $E_{\rm k}=E-m$; the rigidity, $R=pc/Ze$, is also used later on in the paper.
@@ 2.2 Specifics of EC decay in GCRs
The modification of the transport Eq. (1) to account for EC decay, along with the explicit solutions for the LBM and 1D-DM are detailed in B.
@@ 2.3 Dominant processes and timescales
We show in this section the timescales of the processes involved in 1D-DM, consisting of a thin disc of half-width $h$ (source and gas) and a diffusive halo of half-width $L$ (in an infinite plan). The interplay between these timescales shapes the isotopic and elemental GCR fluxes. The escape time from the diffusive volume (or diffusion time) is $t_{\rm diff}=L^{2}/(2D)$. The inelastic, attachment and stripping times in the gaseous disc are obtained from the inverse of the rates defined in Eqs. (2) and (5); the decay time is the inverse of the decay rate Eq. (3).
Table 1: Typical energy and CR charge (or atomic number) dependence of timescales in the 1D-DM (see B) in the relativistic regime ($\beta\approx 1$): diffusion (diff), inelastic interactions (inel), electron radiative attachment (ra) and non-radiative attachment (nra), electron stripping (str), and $\beta$ decay; see A for a full discussion of the attachment and stripping cross-sections.
Rates
$t_{\rm diff}$ $\propto E^{-0.5}$
$t_{\rm inel}$ $\propto A^{-2/3}$
$t_{\rm att}$ $\propto E\,Z^{-5}$ (ra)
$\propto E\,Z^{-5}$ (nra)
$t_{\rm str}$ $\propto Z^{2}/\ln E$
$t_{\rm EC}$ $\propto E$
Figure 1: Timescales for the processes listed in Table 1, as a function of the atomic number $Z$, for three energies (from left to right). The attachment time (pink dotted line) is broken down into the radiative (thin-dotted grey line) and non-radiative (thick-dotted grey line) contributions, see A.1. The black dots represent the effective decay time ($\gamma\,\tau_{\rm EC}$) for each EC-unstable isotopes reported in Table LABEL:tab:EC_clocks (highlighted species are compared to CR data in Sect. 4).
@@
We gather in Table 1 the most salient energy or atomic number dependence of these timescales. In Fig. 1, we show them as a function of the atomic number $Z$, for three energies. We choose the minimum energy 10 MeV/n (left panel) as it corresponds to the lowest IS energies measured by CR experiments (see Sect. 4), and 1 GeV/n as the maximum energy (right panel), because EC decay is blocked above [^2] In passing, this is a striking difference between EC and spontaneous decay, as the former is blocked when $t_{\rm att}\gg\min(t_{\rm diff},\,t_{\rm inel})$, whereas the latter is blocked when $\gamma t_{1/2}\gg t_{\rm diff}$.. As already reported in the literature (e.g., Derome et al., 2019), below a few GeV/n, destruction (blue dashed line) is significant over diffusion (green solid line) for heavy nuclei (i.e., $t_{\rm inel}<t_{\rm diff}$ for $Z\gtrsim 10$). Actually, as discussed in Vecchi et al. (2022), diffusion prevails at increasingly larger energies for growing CR masses.
@@ 3 Impact of EC-decay, uncertainties, and validity
3.1 Preamble
For this analysis, we neglect energy losses, which allows us to use the simple analytical solutions reported in B. Energy losses shape GCR spectra below hundreds of MeV/n, roughly shifting downwards their energy: GCRs at $E_{\rm k/n}$ actually come from $E^{\prime}_{\rm k/n}>E_{\rm k/n}$, that is, an energy where EC decay is less efficient and impacting (see previous section). Consequently, the no-energy loss calculation overestimates the impact of EC decay. We explore below several aspects of EC decay: validity of some approximations, impact of uncertainties, detectability of EC decay in current CR data (Sect. 4). We will show that our conclusions, obtained in the no-energy loss framework, are conservative, that is, they would be even more valid if energy losses were to be accounted for.
@@ 3.2 Impact of EC decay on isotopic IS fluxes
Figure 2: Colour-coded EC-decayed fraction, Eq. (6), in the $t_{1/2}$$Z$ plan for the 1D-DM (top) and LBM (middle), for three IS energies. The two near-vertical lines separate short-lived, intermediate-lived, and long-lived isotopes (see text for details). Symbols indicate existing EC-unstable isotopes (see Table LABEL:tab:EC_clocks), highlighting the names of those for which CR data exist (see Sect. 4).
@@
as a function of a generic $t_{1/2}$ and $Z$. In the above equation, $N_{2\penalty 10000\ {\rm lev}}$, defined in (19), corresponds to the calculation tracking only two states (fully ionised, and single electron attached), as used in the literature, whereas $N_{1\penalty 10000\ {\rm lev}}$ is the calculation without EC decay (only the fully ionised state is enabled).
@@
The near-vertical line on the right-hand side of all panels corresponds to $\gamma\,\tau_{\rm EC}=10\,t_{\rm diff}$: isotopes on the right-hand side of this line mostly escape from the Galaxy before having time to decay, so that their decayed fraction $f<1\%$. The other near-vertical line (on the left) corresponds to the condition $\gamma\tau_{\rm EC}=0.1\,t_{\rm str}$: isotopes on the left-hand side of this line (i.e., where most EC-decay isotopes lie) decay immediately after they attach an electron, hence, their decay is solely controlled by the attachment time. As a result, the decayed fraction is $f\simeq 100\%$ for all species at low energy (left panel), whereas at higher energy (central and right panel), it goes from a few percent up to 100% for growing $Z$, following the behaviour of the attachment time. In the in-between region (i.e., intermediate-lived isotopes), the decay impact depends on the interplay between the EC decay, stripping, and attachment times, and can go from 100% to less than 1%, depending on the energy and the charge of the isotope.
@@
We use the above relation in our calculations, in order to have meaningful comparisons between the two models. However, the 1D-DM is not reducible to the homogeneous LBM for $\beta$-unstable species (Prishchep and Ptuskin, 1975), and we expect the same irreducibility for EC-unstable species. Indeed, this non-equivalence is directly seen from the different formal solutions, Eqs. (31) for the LBM and (32) for the 1D-DM, where the LBM solution is more dominated by the decay term than is the 1D-DM one.
@@
For long-lived isotopes (close or beyond the vertical line on the right), the ratio is $1$, as expected for stable species. For intermediate-lived species with $Z>40$ and $E_{\rm k/n}<100$ MeV/n, the LBM is a factor 1.5 larger than the 1D-DM calculation, whereas just at the transition between short-lived and intermediate-lived species (near-vertical line on the left-hand side of the plot), the LBM can be half the value of the 1D-DM, or even less for $Z<30$ at very low energy (bottom left corner of the left panel). Among the currently measured CR fluxes, $^{44}{\rm Ti}$ is the EC-stable species most impacted by the model choice.
@@ 3.3 Impact of $\sigma^{\rm att}$ and $\sigma^{\rm str}$ uncertainties
In Fig. 4, this ratio is shown for a halved $\sigma^{\rm att}$ (top) and a doubled $\sigma^{\rm str}$ (bottom). This choice allows us to show the two cases with the same colour scale, easing a visual comparison and discussion (of the impact of a 50% uncertainty on these cross-sections). Indeed, both changes lead to less decay, hence a larger $N^{\sigma_{\rm new}}$ and ${\cal R}^{\sigma_{\rm new}}>1$, as it takes a longer (resp. shorter) time to attach (resp. strip) an electron, slowing down and decreasing the decay probability. We stress that if, instead, we were to double $\sigma^{\rm att}$ and half $\sigma^{\rm str}$, this would lead to the opposite behaviour – more decay, smaller $N^{\sigma_{\rm new}}$, and ${\cal R}^{\sigma_{\rm new}}<1$ –, while following a similar pattern in terms of the dependence with $E_{\rm k/n}$, $t_{1/2}$ and $Z$.
@@
In the top panel, we observe a growing impact of EC decay towards lower energies (i.e., from right to left panels), and going from intermediate- to short-lived isotopes for larger $Z$ (upper-left corners of the plots). Long-lived isotopes are insensitive to the modified cross-section, i.e., ${\cal R}^{0.5\sigma^{\rm att}}\!(t_{1/2}\gtrsim 1\,{\rm Myr})\!=\!1$. The impact of halving $\sigma^{\rm att}$ (top panel) is important for short-lived nuclei (${\cal R}^{0.5\sigma^{\rm att}}\!\!=2$), as attachment directly drives the decay time. The region of intermediate-lived nuclei – near the vertical line on the left-hand side – is also impacted, as it is very sensitive to the balance between attachment, stripping, and decay. Looking back at the timescales in Fig. 1, among the species with CR data (i.e., highlighted names in the figure), the curve for $t_{\rm att}$ with a halved decay time would get further away from 53Mn and 41Ca, but closer to 44Ti, 55Fe, etc. This reflects into a maximum impact for the latter, where the decayed flux is changed by more than a factor two at low energy (top left panel in Fig. 4).
@@
A comparison of the bottom and top panels of Fig. 4 shows that changing the stripping and attachment times have very different impacts on the decaying fractions. In addition to have no effect on long-lived isotopes and high energies, there is also no effect on short-lived nuclei, as their decay is not limited by stripping (left side of the plots). Intermediate-lived nuclei are impacted at the same level and for the same nuclei as when halving $\sigma^{\rm att}$ (compare the top and bottom panels), owing to a similar modification of the balance between attachment, stripping, and decay. This impact is ${\cal R}^{2\sigma^{\rm str}}\!\!=2$ at 10 MeV/n (left panel). It decreases at higher energies (right panels), while the most impacted nuclei are shifted towards larger $Z$ and $t_{1/2}$.
@@
    in the bottom-right region of the $t_{1/2}-Z$ plan, we have ${\cal R}^{\sigma^{\rm att}_{\rm new}}\simeq{\cal R}^{\sigma^{\rm str}_{\rm new}}\simeq 1$ and $f\lesssim 1\%$. As there is no foreseeable scenario in which CR data will reach the percent precision on isotopes, large uncertainties on $\sigma^{\rm att}$ and $\sigma^{\rm str}$ (up to a factor two or more) are non-impacting.
@@
    in the top-left region of the $t_{1/2}-Z$ plan, we have ${\cal R}^{\sigma^{\rm att}_{\rm new}}\gtrsim 2$, ${\cal R}^{\sigma^{\rm str}_{\rm new}}\approx 1$ and $f\sim 100\%$. The latter number means that EC-decay decreases the fluxes by a factor $\gtrsim 100$ and is insensitive to uncertainties on $\sigma^{\rm str}$, but that a 50% uncertainty on $\sigma^{\rm att}$ translates to a factor of two uncertainty on this suppression factor: while significant, this does not qualitatively change the conclusions regarding the detectability of EC-decay in CR data (the data precision goes from a few percent to tens of percent, see Sect. 4).
@@
    intermediate-lives nuclei (left-hand side of the near-vertical line) with $Z\approx 30$ (i.e., 44Ti, 55Fe, etc.): this is a region, at low-energy, where both ${\cal R}^{\sigma^{\rm att}_{\rm new}}$ and ${\cal R}^{\sigma^{\rm str}_{\rm new}}$ are $\gtrsim 2$, and $5\lesssim f\lesssim 50\%$ (depending on the energy). Consequently, the cross-section uncertainties maximally impact the conclusions for them: the EC-decay impact can go from very small (few percent) to very significant (up to a factor 2), compared to the CR data precision.
We stress that we obtain similar results for the LBM, and also for multi-level models (not shown). We will come back to these uncertainties in Sect. 4. However, we can already say that, overall, most EC-unstable isotopes are insensitive to large uncertainties on $\sigma^{\rm str}$, while their behaviour is only qualitatively insensitive to uncertainties on $\sigma^{\rm att}$. Only the flux of isotopes $t_{1/2}\approx{\cal O}(\rm kyr)$ at IS energies below a few hundreds of MeV/n is critically sensitive to uncertainties on both cross-sections.

### v2

@@ 2.1 From differential density to modulated fluxes (without EC)
$\begin{split}-\vec{\nabla}\left[D(E)\vec{\nabla}N_{i}-\vec{V}_{c}N_{i}\right]+\frac{\partial}{\partial E}\left[b_{\rm tot}(E)N_{i}-(v/c)^{2}K_{pp}\frac{\partial N_{i}}{\partial E}\right]\\ +\left(\Gamma^{\rm inel}_{i}+\Gamma^{\beta}_{i}\right)N_{i}=q_{i}+\sum_{j>i}\Gamma^{\rm prod}_{j\to i}N_{j}+\sum_{k>i}\Gamma^{\beta}_{k\to i}N_{k}\,.\end{split}$ (1)
In the first line, from left to right, the transport terms are: spatial diffusion $D$ (CR scattering off the magnetic turbulence), convective transport $V_{c}$, energy losses (ionisation, Coulomb, and adiabatic) and gains (momentum diffusion $K_{pp}$). In the second line, the CR destruction (left-hand side) and production from heavier species (right-hand side) involve, respectively, the inelastic cross-section of CR species $i$ on the ISM targets $t$, and the spallation cross-section of $i$ on $t$ to produce $j$:
$\Gamma^{\rm inel}_{i}=\sum_{t\,\in\,\rm ISM}\,n_{t}\,v_{i}\,\sigma_{i+t}^{\rm inel}\quad{\rm and}\quad\displaystyle\Gamma^{\rm prod}_{j\to i}=\sum_{t\,\in\,\rm ISM}\,n_{t}\,v_{i}\,\sigma_{j+t\to i}^{\rm prod}\,.$ (2)
It also involves $\beta$ decay (half-life $t_{1/2}$) as a sink (decay of $i$) or a source (decay of CR $k$ into $i$),
@@
We set $n_{\rm ISM}=\sum_{t}n_{t}=1\,\text{cm}^{-3}$ considering a contribution to the ISM number density from H and He of 90% and 10%, respectively (e.g., Maurin, 2020). Finally, the injection rate $q_{i}$ describes the primary production.
This transport equation can be solved by different approaches and associated codes: numerically after discretisation on spatial and energy coordinates – e.g., DRAGON (Evoli et al., 2018), GALPROP (Porter et al., 2022) and PICARD (Kissmann, 2014) –, via stochastic differential equation based Monte Carlo simulations (e.g., CRPROPA, Merten and Aerdker 2025), or from semi-analytical solutions in simplified geometries based on Green’s functions, series expansion, etc. (e.g., the LBM and 1D-DM in USINE [^1] https://dmaurin.gitlab.io/USINE/, Maurin 2020). Once $N$ is calculated for all CR isotopes, the (nearly isotropic) CR interstellar (IS) flux is given by $\psi^{\rm IS}=N(v/4\pi)$. To compare the modelled fluxes to the measured ones, Solar modulated top-of-atmosphere (TOA) fluxes are calculated from IS ones, using the force-field approximation (Gleeson and Axford, 1967, 1968). The energy is transformed according to
@@
with $Z$ and $A$ the atomic and mass number of the species, $\phi_{\rm FF}$ the solar modulation potential varying from $\sim 500$ MV (Solar minimum) to 1200 MV (Solar maximum). The flux at the top of the atmosphere is related to the interstellar flux by $\psi^{\rm TOA}(E^{\rm TOA})=\psi^{\rm IS}(E^{\rm IS})\times(p^{\rm TOA}/p^{\rm IS})^{2}$. In the above equations, $p$ is the momentum, $E$ is the total energy, and $E_{\rm k/n}=E_{\rm k}/A$ is the kinetic energy per nucleon, with $E_{\rm k}=E-mc^{2}$; the rigidity, $R=pc/Ze$, is also used later on in the paper.
@@ 2.2 Specifics of EC decay in GCRs
The modification of the transport Eq. (1) to account for EC decay, along with the explicit solutions for the LBM and 1D-DM are detailed in Sect. 3.1 and B.
@@ 2.3 Dominant processes and timescales
We show in this section the timescales of the processes involved in 1D-DM, consisting of a thin disc of half-width $h$ (source and gas) and a diffusive halo of half-width $L$ (in an infinite plane). The interplay between these timescales shapes the isotopic and elemental GCR fluxes. The escape time from the diffusive volume (or diffusion time) is $t_{\rm diff}=L^{2}/(2D)$. The inelastic, attachment and stripping times in the gaseous disc are obtained from the inverse of the rates defined in Eqs. (2) and (5); the decay time is the inverse of the decay rate Eq. (3).
Table 1: Typical energy, CR charge (or atomic number) and mass number dependence
of timescales in the 1D-DM (see B) in the relativistic regime ($\beta\approx 1$): diffusion (diff), inelastic interactions (inel), electron radiative attachment (ra) and non-radiative attachment (nra), electron stripping (str), and $\beta$ decay; see A for a full discussion of the attachment and stripping cross-sections. Rates $t_{\rm diff}$ $\propto E^{-0.5}$ $t_{\rm inel}$ $\propto A^{-2/3}$ $t_{\rm att}$ $\propto E\,Z^{-5}$ (ra) $\propto E\,Z^{-5}$ (nra) $t_{\rm str}$ $\propto Z^{2}/\ln E$ $t_{\rm EC}$ $\propto E$
Figure 1: Timescales for the processes listed in Table 1, as a function of the atomic number $Z$, for three energies (from left to right). The attachment time (pink dotted line) is broken down into the radiative (thin-dotted grey line) and non-radiative (thick-dotted grey line) contributions, see A.1. The black dots represent the effective decay time, $t_{\rm EC}=\gamma\tau_{\rm EC}$,
for each EC-unstable isotopes reported in Table LABEL:tab:EC_clocks (highlighted species are compared to CR data in Sect. 4).
@@
We gather in Table 1 the most salient energy, atomic number or mass number dependence of these timescales. In Fig. 1, we show them as a function of the atomic number $Z$, for three energies. We choose the minimum energy 10 MeV/n (left panel) as it corresponds to the lowest IS energies measured by CR experiments (see Sect. 4), and 1 GeV/n as the maximum energy (right panel), because EC decay is blocked above [^2] In passing, this is a striking difference between EC and spontaneous decay, as the former is blocked when $t_{\rm att}\gg\min(t_{\rm diff},\,t_{\rm inel})$, whereas the latter is blocked when $\gamma t_{1/2}\gg t_{\rm diff}$.. As already reported in the literature (e.g., Derome et al., 2019), below a few GeV/n, destruction via inelastic interactions (blue dashed line) is significant over diffusion (green solid line) for heavy nuclei (i.e., $t_{\rm inel}<t_{\rm diff}$ for $Z\gtrsim 10$). Actually, as discussed in Vecchi et al. (2022), diffusion prevails at increasingly larger energies for growing CR masses.
@@ 3 Impact of EC-decay, uncertainties, and validity
3.1 Multi-level transport equation
and
The analytical solutions of the transport equation, for stable and $\beta$-radioactive GCR ions, are gathered in Maurin (2020), for the LBM and 1D-DM. We note that the source term cancels out in all IS flux ratios computed in Section 3 and Section 4; consequently, it is not included in the model. In this work, we extend these solutions to the EC case, without energy redistribution and without galactic wind (for the 1D-DM), which allows us to use the simple analytical solutions reported in B.
Energy losses shape GCR spectra below hundreds of MeV/n, roughly shifting downwards their energy: GCRs at $E_{\rm k/n}$ actually come from $E^{\prime}_{\rm k/n}>E_{\rm k/n}$, that is, an energy where EC decay is less efficient and impacting (see previous section). Consequently, the no-energy loss calculation overestimates the impact of EC decay. We explore below several aspects of EC decay: validity of some approximations, impact of uncertainties, detectability of EC decay in current CR data (Sect. 4). We will show that our conclusions, obtained in the no-energy loss framework, are conservative, that is, they would be even more valid if energy losses were to be accounted for.
Compared to the standard solutions, we now have to follow the various ionisation states of the GCR species to include EC decay. Indeed, whereas CRs are expected to be fully ionised at high-enough energy (see Fig. 1 in Sect. 2.3), several ionisation states may co-exist at lower energy (Letaw et al., 1984b, 1985). As a result, instead of solving the transport equation for the differential density of the CR ion $N$, we have to solve it for a tower of ionised states, denoted below $N_{0}$ (fully ionised), $N_{1}$ (one $e^{-}$ attached), $N_{2}$ (two $e^{-}$ attached), etc. Each level is connected to the one below and above via the electron attachment and stripping rates, $\Gamma^{\rm att}$and $\Gamma^{\rm str}$, the decay rate $\Gamma^{\rm EC}=1/(\gamma\tau_{EC})$ being enabled for all $N_{j>0}$ (i.e., at least one electron attached). The total measured flux of this CR species is then given by
$N_{(n+1)\penalty 10000\ {\rm lev}}=N_{0}+N_{1}+N_{2}+\dots+N_{n}\,$, (8)
@@ 3.2 Impact of EC decay on isotopic IS fluxes
Figure 2: Colour-coded EC-decayed fraction, Eq. (9), in the $t_{\rm EC}$$Z$ plane for the 1D-DM (top) and LBM (bottom), for three IS energies. The two near-vertical lines separate short-lived ($t_{\rm EC}<0.1\,t_{\rm str}$), intermediate-lived ($0.1\,t_{\rm str}<t_{\rm EC}<10\,t_{\rm diff}$), and long-lived isotopes ($t_{\rm EC}>10\,t_{\rm diff}$), where $t_{\rm EC}=\gamma\tau_{\rm EC}$
. Symbols indicate known EC-unstable isotopes (see Table LABEL:tab:EC_clocks), highlighting the names of those for which CR data exist (see Sect. 4).
@@
as a function of a generic $t_{\rm EC}$ and $Z$. In the above equation, $N_{2\penalty 10000\ {\rm lev}}$, defined in (8), corresponds to the calculation tracking only two states (fully ionised, and single electron attached), as used in the literature, whereas $N_{1\penalty 10000\ {\rm lev}}$ is the calculation without EC decay (only the fully ionised state is enabled).
@@
The near-vertical line on the right-hand side of all panels corresponds to $t_{\rm EC}(=\gamma\tau_{\rm EC})=10\,t_{\rm diff}$: isotopes on the right-hand side of this line mostly escape from the Galaxy before having time to decay, so that their decayed fraction $f<1\%$. The other near-vertical line (on the left) corresponds to the condition $t_{\rm EC}=0.1\,t_{\rm str}$: isotopes on the left-hand side of this line (i.e., where most EC-decay isotopes lie) decay immediately after they attach an electron, hence, their decay is solely controlled by the attachment time. As a result, the decayed fraction is $f\simeq 100\%$ for all species at low energy (left panel), whereas at higher energy (central and right panel), it goes from a few percent up to 100% for growing $Z$, following the behaviour of the attachment time. In the in-between region (i.e., intermediate-lived isotopes), the decay impact depends on the interplay between the EC decay, stripping, and attachment times, and can go from 100% to less than 1%, depending on the energy and the charge of the isotope.
@@
We use the above relation in our calculations, in order to have meaningful comparisons between the two models. However, the 1D-DM is not reducible to the spatially homogeneous LBM for $\beta$-unstable species (Prishchep and Ptuskin, 1975), and we expect the same irreducibility for EC-unstable species. Indeed, this non-equivalence is directly seen from the different formal solutions, Eqs. (31) for the LBM and (32) for the 1D-DM, where the LBM solution is more dominated by the decay term than is the 1D-DM one.
@@
For long-lived isotopes (close or beyond the vertical line on the right), the ratio is $1$, as expected for stable species. For intermediate-lived species with $Z>40$ and $E_{\rm k/n}<100$ MeV/n, the LBM is a factor 1.5 larger than the 1D-DM calculation, whereas just at the transition between short-lived and intermediate-lived species (near-vertical line on the left-hand side of the plot), the LBM can be half the value of the 1D-DM, or even less for $Z<30$ at very low energy (bottom left corner of the left of Fig 3). Among the currently measured CR fluxes, $^{44}{\rm Ti}$ is the EC-stable species most impacted by the model choice.
@@ 3.3 Impact of $\sigma^{\rm att}$ and $\sigma^{\rm str}$ uncertainties
In Fig. 4, this ratio is shown for a halved $\sigma^{\rm att}$ (top) and a doubled $\sigma^{\rm str}$ (bottom). This choice allows us to show the two cases with the same colour scale, easing a visual comparison and discussion (of the impact of a factor 2 uncertainty on these cross-sections). Indeed, both changes lead to less decay, hence a larger $N^{\sigma_{\rm new}}$ and ${\cal R}^{\sigma_{\rm new}}>1$, as it takes a longer (resp. shorter) time to attach (resp. strip) an electron, slowing down and decreasing the decay probability. We stress that if, instead, we were to double $\sigma^{\rm att}$ and half $\sigma^{\rm str}$, this would lead to the opposite behaviour – more decay, smaller $N^{\sigma_{\rm new}}$, and ${\cal R}^{\sigma_{\rm new}}<1$ –, while following a similar pattern in terms of the dependence with $E_{\rm k/n}$, $t_{\rm EC}$ and $Z$.
@@
In the top panel, we observe a growing impact of EC decay towards lower energies (i.e., from right to left panels), and going from intermediate- to short-lived isotopes for larger $Z$ (upper-left corners of the plots). Long-lived isotopes are insensitive to the modified cross-section, i.e., ${\cal R}^{0.5\sigma^{\rm att}}\!(t_{\rm EC}\gtrsim 1\,{\rm Myr})\!=\!1$. The impact of halving $\sigma^{\rm att}$ (top panel) is important for short-lived nuclei (${\cal R}^{0.5\sigma^{\rm att}}\!\!=2$), as attachment directly drives the decay time. The region of intermediate-lived nuclei – near the vertical line on the left-hand side – is also impacted, as it is very sensitive to the balance between attachment, stripping, and decay. Looking back at the timescales in Fig. 1, among the species with CR data (i.e., highlighted names in the figure), the curve for $t_{\rm att}$ with a halved decay time would get further away from 53Mn and 41Ca, but closer to 44Ti, 55Fe, etc. This implies a maximum impact for the latter, where the decayed flux is changed by more than a factor two at low energy (top left panel in Fig. 4).
@@
A comparison of the bottom and top panels of Fig. 4 shows that changing the stripping and attachment times have very different impacts on the decaying fractions. In addition to have no effect on long-lived isotopes and high energies, there is also no effect on short-lived nuclei, as their decay is not limited by stripping (left side of the plots). Intermediate-lived nuclei are impacted at the same level and for the same nuclei as when halving $\sigma^{\rm att}$ (compare the top and bottom panels), owing to a similar modification of the balance between attachment, stripping, and decay. This leads to ${\cal R}^{2\sigma^{\rm str}}\!\!=2$ at 10 MeV/n (left panel). It decreases at higher energies (right panels), while the most impacted nuclei are shifted towards larger $Z$ and $t_{\rm EC}$.
@@
    in the bottom-right region of the $t_{\rm EC}$$Z$ plane, we have ${\cal R}^{\sigma^{\rm att}_{\rm new}}\simeq{\cal R}^{\sigma^{\rm str}_{\rm new}}\simeq 1$ and $f\lesssim 1\%$. As there is no foreseeable scenario in which CR data will reach the percent precision on isotopes, large uncertainties on $\sigma^{\rm att}$ and $\sigma^{\rm str}$ (up to a factor two or more) are non-impacting.
@@
    in the top-left region of the $t_{\rm EC}$$Z$ plane, we have ${\cal R}^{\sigma^{\rm att}_{\rm new}}\gtrsim 2$, ${\cal R}^{\sigma^{\rm str}_{\rm new}}\approx 1$ and $f\sim 100\%$. The latter number means that EC-decay decreases the fluxes by a factor $\gtrsim 100$ and is insensitive to uncertainties on $\sigma^{\rm str}$, but that a factor 2 uncertainty on $\sigma^{\rm att}$ translates to a factor of two uncertainty on this suppression factor: while significant, this does not qualitatively change the conclusions regarding the detectability of EC-decay in CR data (the data precision goes from a few percent to tens of percent, see Sect. 4).
@@
    intermediate-lived nuclei (inner region between the two near-vertical lines) with $Z\approx 30$ (i.e., 44Ti, 55Fe, etc.): this is a region, at low-energy, where both ${\cal R}^{\sigma^{\rm att}_{\rm new}}$ and ${\cal R}^{\sigma^{\rm str}_{\rm new}}$ are $\gtrsim 2$, and $5\lesssim f\lesssim 50\%$ (depending on the energy). Consequently, the cross-section uncertainties maximally impact the conclusions for this group of nuclei: the EC-decay impact can go from very small (few percent) to very significant (up to a factor 2), compared to the CR data precision.
We stress that we obtain similar results for the LBM, and also for multi-level models (not shown). We will come back to these uncertainties in Sect. 4. However, we can already say that, overall, most EC-unstable isotopes are insensitive to large uncertainties on $\sigma^{\rm str}$, while they show no qualitative change due to uncertainties on $\sigma^{\rm att}$. Only the flux of isotopes with $t_{\rm EC}\approx{\cal O}(\rm kyr)$ at IS energies below a few hundreds of MeV/n is critically sensitive to uncertainties on both cross-sections.